# The MORRIS Experiment: Magnetic Levitation as a New Probe of Non-Newtonian Gravity

### v1

@@ Introduction
From this potential, we may compute the total force using $\bm{F}(\bm{r})=-\bm{\nabla}U(r)$, yielding
$\bm{F}(\bm{r})=-\frac{G_{\infty}m_{s}m_{p}}{r^{2}}\left[1+\alpha\left(1+\frac{r}{\lambda}\right)e^{-r/\lambda}\right]\hat{\bm{r}}\,.$ (2)
@@
In our experiment, we generate this force via a time-periodic source. We achieve this by using either a spinning disk with three cylindrical masses removed in the case of our short- and medium-term setups, or a driven magnetically levitated particle within a second trap for our long-term stage. We illustrate these setups in Fig. 1. We place the source at a variable central displacement from the levitated particle $r_{0}$, representing the distance between the particle and the midpoint of the oscillation, which has amplitude $R$. The most relevant physical scale is the closest approach distance between a source mass and the particle, $a(r_{0})\equiv r_{0}-R$. By varying this, we obtain sensitivity to violations of the ISL.
@@
At each central position $r_{0}$, we record the motion of the particle along a fixed sensitivity axis $\bm{\zeta}$ using a superconducting pickup coil. The pertinent, time-dependent total force is then
$F_{\mathrm{tot}}(t,r_{0})\equiv\sum_{i=1}^{N_{\mathrm{mass}}}\bm{F}\textbf{(}\bm{r}_{i}(t,r_{0})\textbf{)}\cdot\hat{\bm{\zeta}}\,,$ (3)
@@
where the sum runs over the number of source masses $N_{\mathrm{mass}}$ (three for our short- and medium-term setups and one for our long-term setup). We take $\bm{\zeta}$ to be aligned with the $x$-axis of our experiment, along which we also place the oscillating source.
@@
Figure 1: A: Top view of the proposed short- and medium-term experimental setups for MORRIS, showing the mass-wheel and the trapped particle. The quantities shown are the separation between the driving and test masses at closest approach $a$, the distance modulation $R$, the frequency of the drive $\omega_{d}$, and the total force experienced by the particle $F_{\mathrm{tot}}(t)$. B: Side view of the proposed long-term setup, consisting of two traps. In the driving trap, an AC electromagnetic drive excites the trapped particle. C: The force experienced by the sensing particle with signal period $T_{s}$. The approximate dependence of this force to $a$ and $R$ is explained in Ref. [34].
@@
To simplify our analysis, we decompose the total force into the sum of a constant component, which is irrelevant for our study, and an oscillatory component, $F_{\mathrm{osc}}(t,r_{0})\equiv\mathop{{}\mathcal{F}_{\mathrm{osc}}(r_{0})}\sin(\omega_{s}t)$. The angular frequency $\omega_{s}$ is that of our periodic source, given by $\omega_{s}=N_{\mathrm{mass}}\,\omega_{d}$, with $\omega_{d}$ the angular frequency of the drive. The amplitude $\mathcal{F}_{\mathrm{osc}}(r_{0})$ is given by half the difference between the maximal and minimal driving forces. We visualize the driving force in Fig. 1 and validate our approximation in Ref. [34].
@@
Ultimately, we conduct our analysis in Fourier space by analyzing the force power spectral density (PSD), $\mathcal{P}(\omega)$. Our signal manifests as a monochromatic peak at the angular frequency $\omega_{s}$, with a peak height proportional to $\mathcal{F}_{\mathrm{osc}}^{2}(r_{0})$ [34]. The quantity of interest for our inferencing is the excess power at this frequency normalized by the noise PSD $S_{FF}^{\mathrm{noise}}(\omega_{s})$. The normalized excess power at the central displacement $r_{0}$ after observing for a time $T_{\mathrm{obs}}$ is given by the square of the dimensionless quantity
@@
For the first benchmark point, the fifth force is not effectively screened, resulting in a signal that does not observably violate the ISL over the scale of our experiment. Instead, it appears as a scaling of the gravitational signal. For the remaining benchmark points, ISL violations are expected, with the latter featuring a repulsive force that can lead to a total cancellation of the gravitational signal; this can be seen at $a\approx 15\,\mathrm{mm}$. Since systematic uncertainties can scale our expected signal but are unlikely to mimic the spectral features induced by ISL violations, we are most sensitive to screening lengths of the order of our shortest experiment scale, $\lambda\sim a_{\mathrm{min}}\sim\mathcal{O}(1\,\mathrm{mm}\text{--}10\,\mathrm{mm})$.
Figure 2: The force power spectral density of our levitated particle taken at the resonance frequency of the particle, $\mathcal{P}(\omega_{s})$, with varying closest approach distance $a$. We assume our short-term configuration (c.f. Table 1). Shown is the expected gravitational signal ($\alpha=0)$ with five equally spaced data points (red markers). The fifth-force signals from three benchmark points are also shown, labeled according to $(\alpha,\lambda)$. The noise floor is assumed to be thermal, $S_{FF}^{\mathrm{noise}}(\omega_{s})=S_{FF}^{\mathrm{therm}}$, and is given in Eq. 5.
@@
Experiment We show our proposed experiment, MORRIS, in Fig. 1. At its core is a composite particle consisting of a magnetic core and a non-magnetic weight that is levitated within a superconducting trap and driven by an external periodic force. We will use $\mathrm{Nd_{2}Fe_{14}B}$ magnets due to their high remnant magnetization to aid in signal detection, silica to add mass without increasing eddy-current damping, and lead traps for their high type-I transition temperature (7.2 K). The magnetic trapping mechanism yields high quality factors $Q$ (in excess of $10^{7}$) at low resonance frequencies $f_{0}$ ($\mathord{\sim}10\,\mathrm{Hz}$[35]. Low-frequency operation enables the generation of a time-varying force gradient with a high modulation depth, leading to an increased signal. The gradient is sourced near the resonance frequency of the trapped particle such that the motional amplitude of the particle steadily increases. Our experiment will feature short-, medium-, and long-term stages that will allow us to search for non-Newtonian gravity with increasing sensitivities. They are expected to take one, two, and five years to construct, respectively. We summarize them in Table 1, providing technical details in Ref. [34].
@@
$T_{\mathrm{obs}}~{}\mathrm{[h]}$ $20$ $1000$ $1000$
$T~{}\mathrm{[K]}$ $4$ $1$ $0.1$
@@
In the short- and medium-term setups, the periodic force gradient is sourced by a high-density solid disk of radius $1.25\,\mathrm{cm}$ with three cylindrical masses removed. This density modulation sources a gradient when the disk is spun, such that the driving force varies with a frequency matching the resonance frequency of our particle. We will use tungsten for its high density, ease of machining, and commercial availability. Our proposed long-term realization instead utilizes a direct drive of a particle suspended within a second trap, which then sources a force gradient to the sensing mass. The second particle is driven either through an electrostatic interaction or via magnetic coils. This setup is similar to that proposed for precision measurements of the gravitational constant and in earlier work performed by the group [36, 32]. The geometry of this configuration allows us to achieve a smaller closest approach distance, halving $a_{\mathrm{min}}$, while still guaranteeing adequate shielding from external electromagnetic forces. To minimize this distance, we will reduce the size of the traps from $\mathord{\sim}1\,\mathrm{cm}$, as used in previous work [23], down to $2.5\,\mathrm{mm}$ along the minor axis for the long-term stage.
During the operation of the experiment, it is important that the particle is sufficiently shielded from external disturbances. This includes electromagnetic shielding and vibration isolation. Such shielding has been shown to sufficiently isolate a $0.4\,\mathrm{mg}$ mass to a translational mode temperature of below $4\,\mathrm{K}$ [23], placing the excess noise below the thermal noise for our short-term stage. Higher levels of vibration isolation have been reported by the LIGO collaboration and implementing similar levels of isolation would allow for thermally limited operation down to $100\,\mathrm{mK}$ [37]. Fiducializing to our short-term setup and using $\gamma\equiv\omega_{0}/Q$ as our mechanical dissipation rate, our expected thermal noise is given by
@@
A key advantage of our experimental strategy is that we have excellent control over our source. Not only can we generate a driving force at a well-characterized frequency, such that we can target the signal in Fourier space, but we can also instrument the source masses to have precisely known values and vary the separation between the source and test particle. Placing the source into the vacuum of the setup greatly reduces the minimum separation between the test and source masses, giving us sensitivity to smaller screening lengths $\lambda$.
Moreover, measuring the force at varying separations gives us sensitivity to departures from the ISL. We will therefore make five equally spaced measurements, beginning from the minimum allowed central displacement of each setup $r_{0}^{\mathrm{min}}$: $19.5\,\mathrm{mm}$ for our short- and medium-term stages and $3\,\mathrm{mm}$ for our long-term goal. We will displace the source $2\,\mathrm{mm}$ at a time in the first two setups, improving this to $1\,\mathrm{mm}$ in the long term. Remaining close to the minimal displacements gives us larger signal-to-noise ratios and offers us greater sensitivities. Five data points are sufficient to give us competitive sensitivities when spectral information is available while maintaining a realistic total expedition time of $\mathord{\sim}5T_{\mathrm{obs}}$. From an experimental perspective, sampling at different separations allows for the rejection of systematic errors in both the closest separation $a$ and the force modulation depth by matching the curvature of the results.
@@
Our expected signal is approximately a monochromatic peak in the force PSD [34]. The likelihood function governing the noise-normalized excess power is a non-central $\chi^{2}$ function, $\chi^{2}_{\mathrm{nc}}$, with $k=2$ degrees of freedom [38]. For the $i^{\mathrm{th}}$ data point $p_{i}$, taken at the central displacement $r_{0,i}$, the non-centrality parameter is given by $\kappa^{2}_{i}\equiv\kappa^{2}(r_{0,i})$. To account for potential systematic uncertainties in both the signal and the noise, we include a nuisance pull parameter $\xi$ in our analysis that we ultimately profile out. This enters as a scaling of the non-centrality parameter, such that $\kappa_{i}\rightarrow(1+\xi)\kappa_{i}$. We take $\xi\sim\mathcal{N}(0,\sigma_{\xi}^{2})$, where $\sigma_{\xi}$ is its uncertainty.
@@
Since each data point is identically and independently distributed, the total likelihood is the product of the likelihood of each individual data point. Including the penalty term for the pull parameter, our total likelihood given the observed data vector $\bm{p}\equiv\{p_{i}\}$ is
$\begin{split}\mathcal{L}(\alpha,\lambda,\xi;\bm{p})&\equiv\prod_{i=1}^{N_{\mathrm{data}}}\mathop{{}\chi^{2}_{\mathrm{nc}}\left(p_{i};(1+\xi)^{2}\kappa_{i}^{2},k=2\right)}\\ &\times\frac{1}{\sqrt{2\pi}\sigma_{\xi}}\exp\left(-\frac{\xi^{2}}{2\sigma_{\xi}^{2}}\right)\,,\end{split}$ (6)
@@
where $N_{\mathrm{data}}=5$ is the number of data points we measure at the varying central displacements. Using this likelihood, we define the profile likelihood ratio $\lambda_{\mathrm{PLR}}(\alpha,\lambda=\lambda_{0};\bm{p})$, which profiles away the pull parameter. We then construct the two-sided test statistic $t_{\alpha}\equiv-2\ln\left[\mathop{{}\lambda_{\mathrm{PLR}}(\alpha,\lambda=\lambda_{0};\bm{p})}\right]$ and use it to draw our projected sensitivities [39]. We detail our complete inference pipeline in Ref. [34], which is generally applicable to searches for well-characterized sinusoidal driving forces.
@@
To calculate our projections, we simulate $10^{3}$ pseudo-datasets according to a gravity-only signal. For each $\lambda=\lambda_{0}$, we scan over $\alpha$ to find our limit, which occurs when $t_{\alpha}\approx 3.84$ at the $95\%$ confidence level (CL) [34]. This procedure gives us a distribution of limits, the median of which we take as our best estimator of the limit. We also use these distributions to define $1\sigma$ bands within which we expect each limit to lie. Since our projections for negative $\alpha$ are similar to those for positive values, we choose only to visualize our positive sensitivities, providing the full limit landscape in Ref. [34].
@@
We show our projected $95\%$ CL limits in Fig. 3. All limits exhibit the same characteristic behaviors with varying screening length $\lambda$. We achieve optimal sensitivity when $\lambda\sim a_{\mathrm{min}}$, the minimum closest approach distance of a source mass. This is because the amplitude $\mathcal{F}_{\mathrm{osc}}$ is highly separation dependent, giving the data point at the closest separation the largest signal-to-noise ratio. This results in a turnover at $\lambda\sim 7\,\mathrm{mm}$ for the short- and medium-term realizations and at $\lambda\sim 2\,\mathrm{mm}$ for the long-term stage (a similar scale to this was found in Ref. [32]). Since our projections are based on a generic Yukawa parametrization, they can be readily recast to concrete models predicting the existence of fifth forces, such as novel light scalar or vector mediators [3].
@@
Figure 3: The $95\%$ confidence level projected limits on positive values of $\alpha$ with scale $\lambda$ from MORRIS. We show the limits from our short-, medium-, and long-term configurations, summarized in Table 1. The bands around each limit illustrate their $1\sigma$ range. The solid regions indicate the existing fifth-force limits from laboratory [40, 12, 13, 41] and geophysical [3] experiments. The transparently shaded regions show the theoretically preferred regions for light moduli stemming from supersymmetric string theory [8, 9].
We compare our projections to existing laboratory and geophysical constraints in the $\lambda$ range of interest. The former consists of torsion balance [40, 12] and pendulum [13, 41] experiments, while the latter includes precision measurements of Earth’s gravitational acceleration [3]. Our short-term sensitivity, which is based on an immediately realizable configuration, provides a proof-of-concept for our experiment and is sub-leading compared to current constraints. However, following a set of simple upgrades, a medium-term experiment will have leading sensitivity, constraining $\alpha\lesssim 10^{-4}$ at a screening length of $\lambda\approx 4\,\mathrm{mm}$. Our long-term experiment is sensitive to new parameter space over the screening length range $0.3\,\mathrm{mm}$ to $10\,\mathrm{mm}$, constraining $\alpha\lesssim 10^{-5}$ at $\lambda\approx 2\,\mathrm{mm}$ and surpassing laboratory limits by two orders of magnitude.

### v2

# Magnetic Levitation as a New Probe of Non-Newtonian Gravity

@@ Introduction
From this potential, we may compute the total force using $\boldsymbol{F}(\boldsymbol{r})=-\boldsymbol{\nabla}U(r)$:
$\boldsymbol{F}(\boldsymbol{r})=-\frac{G_{\infty}m_{s}m_{p}}{r^{2}}\left[1+\alpha\left(1+\frac{r}{\lambda}\right)e^{-r/\lambda}\right]\hat{\boldsymbol{r}}\,$. (2)
@@
In our experiment, we generate this force via a time-periodic source. We achieve this by using either a spinning disk with three cylindrical masses removed in the case of our short- and medium-term setups, or a driven magnetically levitated particle within a second trap for our long-term stage. We illustrate these setups in Fig. 1. We place the source at a variable central displacement from the levitated particle $r_{0}$, representing the distance between the particle and the midpoint of the oscillation, possessing amplitude $R$. The most relevant physical scale is the closest approach distance between a source mass and the particle, $a(r_{0})\equiv r_{0}-R$. By varying this, we obtain sensitivity to ISL violations.
@@
At each central position $r_{0}$, we record the motion of the particle along a fixed sensitivity axis $\boldsymbol{\zeta}$ using a superconducting pickup coil coupled to a SQUID. The time-dependent total force is then
$F_{\mathrm{tot}}(t,r_{0})\equiv\sum_{i=1}^{N_{\mathrm{mass}}}\boldsymbol{F}\textbf{(}\boldsymbol{r}_{i}(t,r_{0})\textbf{)}\cdot\hat{\boldsymbol{\zeta}}\,$, (3)
@@
where the sum runs over the number of source masses $N_{\mathrm{mass}}$. We align $\boldsymbol{\zeta}$ with the $x$-axis of our experiment, along which we also place the oscillating source.
@@
Figure 1: A: Top view of the proposed short- and medium-term experimental setups for MORRIS, showing the mass-wheel and the trapped particle. We include the separation between the driving and test masses at closest approach $a$, the distance modulation $R$, the frequency of the drive $\omega_{d}$, and the total force experienced by the particle $F_{\mathrm{tot}}(t)$. B: Side view of the proposed long-term setup, consisting of two traps. In the driving trap, an AC electromagnetic drive excites the trapped particle. C: The force experienced by the sensing particle with signal period $T_{s}$. The dependence of this force to $a$ and $R$ is explained in Ref. [1].
@@
To simplify our analysis, we decompose the total force into the sum of a constant component, which is irrelevant for our study, and an oscillatory component, $F_{\mathrm{osc}}(t,r_{0})\equiv\mathop{{}\mathcal{F}_{\mathrm{osc}}(r_{0})}\sin(\omega_{s}t)$. The angular frequency $\omega_{s}=N_{\mathrm{mass}}\omega_{d}$ is that of our periodic source, with $\omega_{d}$ the angular frequency of the drive. The amplitude $\mathcal{F}_{\mathrm{osc}}(r_{0})$ is given by half the difference between the maximal and minimal driving forces. We visualize the driving force in Fig. 1 and validate our approximation in Ref. [1].
@@
Ultimately, we conduct our analysis in Fourier space by analyzing the force power spectral density (PSD), $\mathcal{P}(\omega)$. Our signal manifests as a monochromatic peak at the angular frequency $\omega_{s}$, with a peak height proportional to $\mathcal{F}_{\mathrm{osc}}^{2}(r_{0})$ [1]. The quantity of interest for our inferencing is the excess power at this frequency normalized by the noise PSD $S_{FF}^{\mathrm{noise}}(\omega_{s})$. The normalized excess power at the central displacement $r_{0}$ after observing for a time $T_{\mathrm{obs}}$ is given by the square of the quantity
@@
For the first benchmark point, the fifth force is not effectively screened, resulting in a signal that does not observably violate the ISL; it instead appears as a scaling of the gravitational signal. For the remaining benchmark points, ISL violations are expected, with the latter featuring a repulsive force that can lead to a cancellation of the gravitational signal. This can be seen at $a\approx 15\,\mathrm{mm}$. Since systematic uncertainties can scale our signal but are unlikely to mimic the spectral features induced by ISL violations, we are most sensitive to screening lengths of the order of our shortest experimental scale, $\lambda\sim a_{\mathrm{min}}\sim\mathcal{O}(1\,\mathrm{mm}\text{--}10\,\mathrm{mm})$.
Figure 2: The force power spectral density of our levitated particle taken at the resonance frequency of the particle, $\mathcal{P}(\omega_{s})$, with varying closest approach distance $a$. We assume our short-term configuration (c.f. Table 1). Shown is the expected gravitational signal ($\alpha=0)$ with five equally spaced data points (red markers). The fifth-force signals from three benchmark points are also shown, labeled according to $(\alpha,\lambda)$. The noise floor is assumed to be thermal, $S_{FF}^{\mathrm{noise}}(\omega_{s})=S_{FF}^{\mathrm{therm}}$ (c.f. Eq. 5).
@@
Experiment We present our proposed experiment, MORRIS, in Fig. 1. At its core is a composite particle consisting of a magnetic core and a non-magnetic weight that is levitated within a superconducting trap and driven by an external periodic force. We will use $\mathrm{Nd_{2}Fe_{14}B}$ magnets due to their high remnant magnetization to aid in signal detection, silica to add mass without increasing eddy-current damping, and lead traps for their high type-I transition temperature (7.2 K). The magnetic trapping mechanism yields high quality factors $Q$ (in excess of $10^{7}$) at low resonance frequencies $f_{0}$ ($\mathord{\sim}10\,\mathrm{Hz}$[40]. Low-frequency operation enables the generation of a time-varying force gradient with a high modulation depth, leading to an increased signal. The gradient is sourced near the resonance frequency of the trapped particle, such that its motional amplitude steadily increases. This will be recorded via a pickup coil coupled to a SQUID—the conversion from SQUID signal to motion will be made either through thermal noise calibration or by injecting an external flux signal [15].
Our experiment features short-, medium-, and long-term stages that will allow us to search for non-Newtonian gravity with increasing sensitivities. They are expected to take one, two, and five years to construct, respectively. We summarize them in Table 1, providing technical details in Ref. [1].
@@
$T_{\mathrm{obs}}~\mathrm{[h]}$ 2020 $1000$ $1000$
$T~\mathrm{[K]}$ 44 11 $0.1$
@@
In the short- and medium-term setups, the force is sourced by a high-density solid disk of radius $1.25\,\mathrm{cm}$ with three cylindrical masses removed. This density modulation sources a gradient when the disk is spun, such that the driving force varies with a frequency matching the resonance frequency of our particle. We will use tungsten for its high density, ease of machining, and commercial availability. Unlike in previous work, we will place the driving wheel inside of the cryogenic cooling system to minimize the shortest approach distance and attain sensitivities down to $\lambda\sim 1\,\mathrm{mm}$ [15].
Our proposed long-term realization instead utilizes a direct drive of a particle suspended within a second trap. The second particle is driven either through an electrostatic interaction or via magnetic coils. This setup is similar to that proposed for precision measurements of the gravitational constant and in earlier work performed by the group [21, 38]. The geometry of this configuration allows us to achieve a smaller closest approach distance, halving $a_{\mathrm{min}}$, while still guaranteeing adequate shielding from external electromagnetic forces. To minimize this distance, we will reduce the size of the traps from $\mathord{\sim}1\,\mathrm{cm}$ down to $2.5\,\mathrm{mm}$ along the minor axis [15].
During the experiment, sufficiently shielding the particle from external disturbances is crucial. This includes electromagnetic shielding and vibration isolation. Such shielding has been shown to adequately isolate a $0.4\,\mathrm{mg}$ mass to a translational mode temperature of below $4\,\mathrm{K}$ [15], placing the excess noise below the thermal noise for our short-term stage. Higher levels of vibration isolation have been reported by the LIGO collaboration and implementing similar levels of isolation would allow for thermally limited operation down to $100\,\mathrm{mK}$ [31]. Moreover, we will optimize the coupling to the particle’s motion such that the SQUID-readout noise is below the thermal noise floor, achieving this by increasing the number of pick-up coil windings and the geometrical coupling to the particles motion. Fiducializing to our short-term setup and using $\gamma\equiv\omega_{0}/Q$ as our mechanical dissipation rate, our expected thermal noise is
@@
A key advantage of our experimental strategy is that we have excellent control over our source. Not only can we generate a driving force at a well-characterized frequency, such that we can target the signal in Fourier space, but we can also instrument the source masses to have precisely known values and vary the separation between the source and test particle. Placing the source in the vacuum greatly reduces the minimum separation between the test and source masses, giving us sensitivity to smaller screening lengths $\lambda$.
Furthermore, measuring the force at varying separations gives us sensitivity to ISL violations. We will therefore make five equally spaced measurements, beginning from the minimum allowed central displacement of each setup $r_{0}^{\mathrm{min}}$: $19.5\,\mathrm{mm}$ for our short- and medium-term stages and $3\,\mathrm{mm}$ for our long-term goal. We will displace the source $2\,\mathrm{mm}$ at a time in the first two setups, improving this to $1\,\mathrm{mm}$ in the long term. Remaining close to the minimal displacements gives us larger signal-to-noise ratios. Five data points are sufficient to give us competitive sensitivities when spectral information is available while maintaining a realistic total expedition time of $\mathord{\sim}5T_{\mathrm{obs}}$. From an experimental perspective, sampling at different separations allows for the rejection of systematic errors in both the closest separation $a$ and the force modulation depth by matching the curvature of the results.
@@
Our expected signal is a monochromatic peak in the force PSD [1]. The likelihood function governing the noise-normalized excess power is a non-central $\chi^{2}$ function, $\chi^{2}_{\mathrm{nc}}$, with $k=2$ degrees of freedom [20]. For the $i^{\mathrm{th}}$ data point $p_{i}$, taken at the central displacement $r_{0,i}$, the non-centrality parameter is $\kappa^{2}_{i}\equiv\kappa^{2}(r_{0,i})$. To account for potential systematic uncertainties in both the signal and the noise, we include a nuisance pull parameter $\xi$ in our analysis that we ultimately profile out. This enters as a scaling of the non-centrality parameter, such that $\kappa_{i}\rightarrow(1+\xi)\kappa_{i}$. We take $\xi\sim\mathcal{N}(0,\sigma_{\xi}^{2})$, where $\sigma_{\xi}$ is its uncertainty.
@@
Since each data point is identically and independently distributed, the total likelihood is the product of the likelihood of each individual data point. Including the penalty term for the pull parameter, our likelihood given the observed data vector $\boldsymbol{p}\equiv\{p_{i}\}$ is
[…]
@@
where $N_{\mathrm{data}}=5$ is the number of measurements. Using this likelihood, we define the profile likelihood ratio $\lambda_{\mathrm{PLR}}(\alpha,\lambda=\lambda_{0};\boldsymbol{p})$, which profiles away the pull parameter. We then construct the two-sided test statistic $t_{\alpha}\equiv-2\ln\left[\mathop{{}\lambda_{\mathrm{PLR}}(\alpha,\lambda=\lambda_{0};\boldsymbol{p})}\right]$ and use it to draw our projected sensitivities [10]. We detail our inference pipeline in Ref. [1], which is generally applicable to searches for well-characterized sinusoidal driving forces.
@@
To calculate our projections, we simulate $10^{3}$ pseudo-datasets according to a gravity-only signal. For each $\lambda=\lambda_{0}$, we scan over $\alpha$ to find our limit, which occurs when $t_{\alpha}\approx 3.84$ at the $95\%$ confidence level (CL) [1]. This procedure gives us a distribution of limits, the median of which we take as our best estimator of the limit. We also use these distributions to define $1\sigma$ limit bands. Since our projections for positive and negative $\alpha$ are similar, we only visualize our positive sensitivities, providing the full limit landscape in Ref. [1].
@@
We show our projected $95\%$ CL limits in Fig. 3. All limits exhibit the same characteristic behaviors with varying screening length $\lambda$. We achieve optimal sensitivity when $\lambda\sim a_{\mathrm{min}}$, the minimum closest approach distance of a source mass. This is because the amplitude $\mathcal{F}_{\mathrm{osc}}$ is highly separation dependent, giving the data point at the closest separation the largest signal-to-noise ratio. This leads to a turnover at $\lambda\sim 7\,\mathrm{mm}$ for the short- and medium-term realizations and at $\lambda\sim 2\,\mathrm{mm}$ for the long-term stage (a similar scale to this was found in Ref. [38]). Since our projections are based on a generic Yukawa parametrization, they can be readily recast to concrete models predicting the existence of fifth forces, such as novel light scalar or vector mediators [3].
@@
Figure 3: The $95\%$ confidence level projected limits on positive values of $\alpha$ with scale $\lambda$ from MORRIS. Shown are the limits from our short-, medium-, and long-term configurations, summarized in Table 1. The bands around each limit illustrate their $1\sigma$ range. The solid regions indicate the existing fifth-force limits from laboratory [24, 27, 44, 37] and geophysical [3] experiments. The transparent regions show the preferred regions for light moduli arising from supersymmetric string theory [12, 13].
We compare our projections to existing laboratory and geophysical constraints in the $\lambda$ range of interest. The former consists of torsion balance [24, 27] and pendulum [44, 37] experiments, while the latter includes precision measurements of Earth’s gravitational acceleration [3]. Our short-term sensitivity, which is based on an immediately realizable configuration, provides a proof-of-concept for our experiment. Following a set of simple upgrades, a medium-term experiment will have leading sensitivity and constrain $\alpha\lesssim 10^{-4}$ at a screening length of $\lambda\approx 4\,\mathrm{mm}$. Our long-term experiment is sensitive to new parameter space over the screening length range $0.3\,\mathrm{mm}$ to $10\,\mathrm{mm}$, constraining $\alpha\lesssim 10^{-5}$ at $\lambda\approx 2\,\mathrm{mm}$ and surpassing laboratory limits by two orders of magnitude.